40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13353     Entry time: Tue Oct 3 01:32:39 2017     In reply to: 13315
Author: gautam 
Type: Update 
Category: LSC 
Subject: Laser intensity noise coupling to MICH (simulated) 

GV Oct 6: This coupling is probably not correct - Finesse outputs TF magnitude in units of W/W, and not W/RIN

Since I was foiled (by lack of DAC) in my attempt to measure the coupling of laser intensity noise to MICH in the DRMI (no arms) configuration, I decided to try understanding the effect with a simulation.

For this purpose, I used my DRMI Finesse model - this had mirror positions tuned for locking and photodiode demod phases tuned to give a sensing matrix model that wasn't too far from an actual measurement (within factor of a few). So the model seems okay for a first pass at estimating this coupling.

Measuring transfer functions in Finesse is straightforward - use the fsig command to modulate some quantity (in this case the input beam intensity), and use the pd2 detector to demodulate the effect of this modulation at the port of interest (in this case AS55_Q).

**Note that to apply a modulation to an input beam (i.e. Laser) in Finesse, the keyword for the "type" argument given to fsig is "amp" and not "amplitude" as the manual would had me believe. In fact, there seem to be quite a few such caveats. The best way to figure this out is to go to the pykat installation directory, find the file components.py, and look for the fsig_name for the component of interest. It is also indicated in the same file, via the canFsig argument, if that property of the component can be modulated for transfer function measurements.  

Attachment #1 shows the result of such a sweep.

To estimate what the actual contribution to the displacement noise is, I used the DQ-ed MC transmission (recorded at 1024Hz) from the DRMI lock, computed the ASD using scipy.signal.welch, divided by the nominal MC transmission of ~15,000 counts to convert to RIN/rtHz. The RIN was then multiplied by the above calculated coupling function, and divided by the sensing matrix element for AS55_Q (in units of W/m) to give the curve shown in Attachment #2. If we believe the simulation, then Laser Intensity Noise shouldn't be the limiting noise between 10Hz-1kHz. 

I will of course measure the actual coupling and see how it lines up with Attachment #1 - would be a nice additional validation of the Finesse model. I will also try using the Finesse model to estimate some other coupling transfer functions (e.g. Laser Frequency Noise, Oscillator Noise).

Quote:

The absence of evidence is not evidence of absence.

 

Attachment 1: MICH_intensityNoiseCoupling.pdf  142 kB  | Hide | Hide all
MICH_intensityNoiseCoupling.pdf
Attachment 2: MICH_intensityNoiseASD.pdf  164 kB  | Hide | Hide all
MICH_intensityNoiseASD.pdf
ELOG V3.1.3-